OpenStudy (anonymous):

Suppose a large soup company wants to minimize the surface area of a can that is to contain 64 cubic inches. What should be the dimensions of the can?

OpenStudy (phi):

This is a calculus problem? First, we assume a can is a cylinder (otherwise, we would choose a sphere, which minimizes surface area per volume) So, what are the formulas for 1) surface area of a cylinder 2) volume of a cylinder (google it if you don't know)

OpenStudy (anonymous):

yes it's calc. 1) SA= [2\pi r ^{2}+2\pi rh\] 2) V=[\pi r ^{2}h\]

OpenStudy (phi):

use eq 2 with V= 64 and solve for h in terms of r substitute this expression for h into the first equation. you now have an expression for surface area as a function of r take the derivative wrt to r, set to 0, and solve for r then find h from the 2nd equation

OpenStudy (anonymous):

how do I solve it when I have a negative exponent?

OpenStudy (phi):

I think the equation is \[ A= 2 \pi r^2 +128 r^{-1} \] Are you asking how to take the derivative? use the power rule \[\frac{d }{dx}x^n= n x^{n-1}\]

OpenStudy (anonymous):

I can do the derivative to solve for the min, but when I do that I get: \[4\pi r-128r ^{-2}=0\] How do I solve for r?

OpenStudy (phi):

multiply both sides by r^2

OpenStudy (anonymous):

mkay. thanks

OpenStudy (phi):

you should get \[ 4\pi r^3 -128= 0\] which gives \[ r= (\frac{32}{\pi} )^{\frac{1}{3}}\]

OpenStudy (anonymous):

yupp that's what I got. Then I just plug it back into where I solved for h in terms of r and get h

OpenStudy (phi):

yes. I got r= 2.1677 and h= 4.3354 (to 4 decimals)

OpenStudy (anonymous):

I got the same thing. Thanks for the help :) much appreciated